# Kolmogorov-Arnold Attention: Is Learnable Attention Better For Vision Transformers?

### v2

@@ B.1.5 Implementing Fourier KArAt for Object Detection and Semantic Segmentation
B.1.6 Plotting Loss Landscapes
@@
B.2 Computation Time, FLOPS, Memory Requirement, and Throughput of Fourier KArAt
The overall computation for Fourier KArAt variants is higher than their conventional softmax MHSA, and we have delineated it in Figure 8. Primarily, the Fourier KArAt variants have a longer training time. We show the training time comparison between the traditional MHSA and its Fourier KArAt versions for 100 epochs on CIFAR-10, CIFAR-100, and ImageNet-1K datasets for all the models (ViT-Tiny, -Small, and -Base) in Figures 8– 8, respectively. We only compare the best-performing Fourier KArAt models ($G_{1}B$ and $G_{1}U$ for ViT-Base, and $G_{3}B$ and $G_{3}U$ for ViT-Tiny & ViT-Small) with their traditional softmax MHSA counterparts. We also observe that universal mode $G_{n}U$ training times are consistently slightly less than the blockwise modes $G_{n}B$.
During the training, we monitored the GPU memory requirements, and as expected, Fourier KArAt variants utilize significantly more memory than traditional MHSA. In particular, the GPU memory requirements scale by 2.53×2.5-3$\times$2.5 - 3 ×, compared to the traditional softmax MHSA.
@@ B.2 Computation Time, FLOPS, Memory Requirement, and Throughput of Fourier KArAt
Figure 8: A detailed comparison of computing requirements.We compare the training times for 100 epochs with the hyperparameter settings given in Table 5 for all the datasets CIFAR-10, CIFAR-100, and ImageNet-1K. We also compare the throughputs of different models on ImageNet-1K; the throughput results will be similar for other datasets as the input size is $224\times 224$.
@@
B.3 Ablation Study with KAN and Fourier KArAt
@@
B.3.1 Experiments with KANs Operated on B-Spline Basis
We test the performance of the B-spline KANs in the classification task on the CIFAR-10 [40], CIFAR-100 [40], MNIST [42, 43], and Fashion-MNIST [77] datasets. We performed extensive experiments to see if there are any benefits to choosing B-Splines for the basis functions in KAN layers. Table 13 shows the Top-1 accuracies from different variants of a Deep KAN. In this set of experiments, we closely followed earlier works [81, 51]. Hyperparameters involved in these experiments include the order of the B-spline $k$, grid size $G$, grid range $[-I,I]$, layers, and width. Although the KANs with B-Spline basis yield high accuracies in the small-scale MNIST and Fashion-MNIST datasets, they fail to generalize over larger datasets (CIFAR-10 and -100).
@@ B.3.1 Experiments with KANs Operated on B-Spline Basis
B.3.2 Ablation with the Hidden Dimension, $r$ in Fourier KArAt
@@
B.3.3 Enforcing an Extreme Low Rank Structure in Fourier KArAt
While our modular design of Fourier KArAt avoids the computational requirement of $\Phi^{i,j}$ by using a low hidden dimension $r$, it also enforces a low-rank structure in the attention matrix ${\cal A}^{i,j}$. In this context, we attempt to find the best possible value for $r$; see §B.3.2. However, we want to investigate how low the hidden dimension $r$ can be used without substantially compromising the performance of learnable Fourier KArAt. The significance of this question lies in the information bottleneck created by an extremely low hidden dimension $r$ that helps to understand the tradeoff between computing requirements and final trained model quality. To this end, we experiment with extreme low-rankness in the hidden dimension, $r=2,4$, and report our findings in Table 17. Although $r=2,4$ are not optimal, and we observe an insignificant drop in the performance across both datasets (CIFAR-10 and CIFAR-100), it comes at a reduced computing requirement, especially required VRAM; see GPU memory in Table 17. For instance, ViT-Tiny+$G_{3}B$ with $r=2$ outperforms the vanilla variant, only has a modest 0.03M parameters increment from the vanilla ViT-Tiny, albeit similar GFLOPs, but an extra 1.07GB of GPU memory usage. Compared to the $r=12$ variant, the memory usage is 3.47 GB lower, a total of 0.73M parameters less, but the relative decrease in accuracy is only 2.96% lower on the CIFAR-10 dataset. In ViT-Base backbones, the performance remains comparable, and even in the case of ViT-Base+$G_{1}B$ on CIFAR-100, the variant with $r=2$ surpasses the original variant with $r=12$. Although there is no concrete strategy, this observation supports our claim on a probable research direction involving parameter reduction (discussed in §6) to improve scalability in such over-parameterized models. This also indicates that for over-parameterized larger models, the scope of parameter reduction may find a model with learnable attention that can outperform its traditional counterpart across multiple datasets.
B.3.4 Ablation on the Impact of grid size, $G$ on Fourier KArAt
@@
Figure 11: Optimization path for ViT-Tiny and ViT-Base (the smallest and the largest model) along the two largest principal component directions of the successive change of model parameters. We show the loss contours along with the trajectory of the optimizers. Perturbed contours indicate a corresponding non-smooth, spiky loss surface.
@@
Table 18: Ablation study on the operator variants in Fourier KArAt. Note ${\widehat{\Phi}(W{\cal A})}$ and ${W\widehat{\Phi}({\cal A})}$ refer to order of using linear projector $W$ and operator $\widehat{\Phi}$, and $\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ refers to two consecutive learnable operators $\widehat{\Phi}_{[1]}$ and $\widehat{\Phi}_{[2]}$.
@@
B.3.5 Does Fourier KArAt Require $\ell_{1}$-Projection?
@@
B.3.6 More Discussion on Blockwise and Universal Configuration
Although the primary inspiration behind the universal configuration comes from the shared basis functions in KAT [80], it aligns with the fact that softmax is a fixed function conventionally being used across all heads of all encoder layers in the ViT architectures. This raises the question of whether the token-to-token interaction in the self-attention can be modeled using a single universal function, or each instance of attention operations needs dynamic modeling using learnable functions. To investigate further, we compared the learned coefficients $\{a_{pqm}$, $b_{pqm}\}$ from Equation (1) and the weight matrix $W$ of the linear projector following the learnable basis, for each head across all the encoder layers in ViT-Tiny+Fourier KArAt. Thus, we obtain 36 sets (3 heads for each of the 12 layers) of such coefficients and weights from the model with blockwise configuration $G_{3}B$, and 3 sets from the model with universal configuration $G_{3}U$. Although mostly centered at zero, the distribution of each of these sets of parameters varies significantly from the others. The difference is mostly visible in the presence of long tails and the length of the tails. This characteristic is present not only in the blockwise configuration, but the three heads of the universal configuration also show a similar pattern.
@@ B.3.6 More Discussion on Blockwise and Universal Configuration
B.4 Alternate Variants of Fourier KArAt
B.4.1 Alternative Approaches to the Lower Rank Attention Structures
In the main paper, we approximated the effect of the operator, ${\Phi^{i,j}}\in\mathbb{R}^{N\times N}$ casting it as the product of two rank-$r$ operators—one operator with learnable activation, $\widehat{\Phi},$ and the other is the learnable the linear transformation matrix, $W$ such that $\Phi=\widehat{\Phi}W$. Here, we abuse the notations for simplicity. However, a natural question is how many different configurations are possible with $\widehat{\Phi}$ and $W$ such that they can approximate the effect of $\Phi.$ Specifically, we use the following configurations — (i) [^4] This configuration was used in the main paper. $W\widehat{\Phi}(\mathcal{A}),$ where $\widehat{\Phi}\in\mathbb{R}^{r\times N}$ and $W\in\mathbb{R}^{N\times r}$, (ii) $\widehat{\Phi}(W{\cal A}),$ where $W\in\mathbb{R}^{r\times N}$ and $\widehat{\Phi}\in\mathbb{R}^{N\times r}$, and (iii) $\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$, where $\widehat{\Phi}_{[1]}\in\mathbb{R}^{r\times N}$ and $\widehat{\Phi}_{[2]}\in\mathbb{R}^{N\times r}$ are low-rank operators for learning activation functions. In the first configuration, $\widehat{\Phi}(\cdot)$ acts on each row of ${\cal A}$ to produce an $r$-dimensional vector, and then $W$ projects it back to an $N$-dimensional subspace. In the second configuration, $W$, first projects each row of ${\cal A}$ to produce an $r$-dimensional vector, and then $\widehat{\Phi}$ with learnable activation produces $N$-dimensional vectors. In the third configuration, $\widehat{\Phi}_{[1]}$ first produces $r$-dimensional vector from rows of ${\cal A}$ by learning activations, and $\widehat{\Phi}_{[2]}$ obtains $N$-dimensional vectors by learning a second level of activations.
B.4.2 Fourier KArAt + softmax Attention —A Hybrid version
B.4.3 More Combinations
B.5 Spectral Analysis of Attention
We use all 3 heads in the last encoder block of ViT-Tiny on 5 randomly sampled images from the CIFAR-10 validation set. There are a total of 15 singular vectors (each of 197 dimensions) for any attention matrix of shape $197\times 197$, where the singular values are arranged in non-increasing order. Let $\sigma_{i}$ be the $i^{\rm th}$ singular value across all heads and samples (it is permutation invariant). For each $i\in[197],$ we plot $[\ln(\sigma_{i}^{\min}),\ln(\bar{\sigma_{i}}),\ln(\sigma_{i}^{\max})]$, where $\bar{\sigma_{i}}$ is the average of $i^{\rm th}$-indexed singular value across all samples and heads.

### v3

@@ C.1.5 Implementing Fourier KArAt for Object Detection and Semantic Segmentation
C.1.6 Plotting Loss Landscapes
@@
C.2 Generalizability: Detection & Segmentation—Continued
We choose Mask RCNN [87]-based framework for this purpose and employ a feature pyramid network [90] based on ViT-Det [41]; it is non-trivial to implement Fourier KArAt, see §C.1.5. We trained ViT-Det [41] using ViT-Base backbone on the MS COCO [42] dataset for 50 epochs with two settings: (i) fine-tuning on the ImageNet-1K pre-trained weights on traditional softmax attention, and (ii) random initialization. Table 12 shows that for the fine-tuning on ImageNet pre-trained weights, Fourier-KArAt shows $\sim 4-13\%$ gap in average precision (AP) from its vanilla variant; their qualitative performance in Figure 8 is similar. Overall, for ViT-Det, Fourier-KArAt performs inferiorly to its conventional counterpart in detection and segmentation. Interestingly, Figure 9 shows, vanilla ViT-Det using softmax activation reaches its peak performance quickly (within 12 epochs), Fourier-KArAt delays in achieving it (within 30 and 45 epochs for $G_{1}B$ and $G_{1}U$, respectively), proving the incompatibility of softmax-based weights initialization. It is, however, still a good initialization for KArAt as the performance is better than random initialization.
Table 12: Fourier KArAt on object detection and instance segmentation tasks on the MS COCO [42] dataset. The header Box and Mask refer to detection and segmentation tasks, respectively.
Model Initialization Box Mask
AP AP50 AP75 AP AP50 AP75
ViT-Det-Base ViT-Base on ImageNet-1K 32.34 49.16 34.43 28.35 46.04 29.48
+ $G_{1}B$ 22.48 36.82 23.13 20.11 34.09 20.46
+ $G_{1}U$ 26.68 43.25 27.88 24.01 40.21 24.66
ViT-Det-Base Random Initialization 15.28 26.88 15.22 13.39 24.34 12.86
+ $G_{1}B$ 10.32 18.77 10.05 8.94 16.51 8.50
+ $G_{1}U$ 10.99 20.04 10.82 9.69 17.99 9.25
C.3 Computation Time, FLOPS, Memory Requirement, and Throughput of Fourier KArAt
The overall computation for Fourier KArAt variants is higher than their conventional softmax MHSA, and we have delineated it in Figure 7. Primarily, the Fourier KArAt variants have a longer training time. We show the training time comparison between the traditional MHSA and its Fourier KArAt versions for 100 epochs on CIFAR-10, CIFAR-100, and ImageNet-1K datasets for all the models (ViT-Tiny, -Small, and -Base) in Figures 7– 7, respectively. We only compare the best-performing Fourier KArAt models ($G_{1}B$ and $G_{1}U$ for ViT-Base, and $G_{3}B$ and $G_{3}U$ for ViT-Tiny & ViT-Small) with their traditional softmax MHSA counterparts. We also observe that universal mode $G_{n}U$ training times are consistently slightly less than the blockwise modes $G_{n}B$.
During the training, we monitored the GPU memory requirements, and as expected, Fourier KArAt variants utilize significantly more memory than traditional MHSA. In particular, the GPU memory requirements scale by $2.5-3\times$, compared to the traditional softmax MHSA.
@@ C.3 Computation Time, FLOPS, Memory Requirement, and Throughput of Fourier KArAt
Figure 7: A detailed comparison of computing requirements.We compare the training times for 100 epochs with the hyperparameter settings given in Table 5 for all the datasets CIFAR-10, CIFAR-100, and ImageNet-1K. We also compare the throughputs of different models on ImageNet-1K; the throughput results will be similar for other datasets, as the input size is $224\times 224$.
@@
C.4 Ablation Study with KAN and Fourier KArAt
@@
C.4.1 Experiments with KANs Operated on B-Spline Basis
We test the performance of the B-spline KANs in the classification task on the CIFAR-10 [35], CIFAR-100 [35], MNIST [91, 86], and Fashion-MNIST [92] datasets. We performed extensive experiments to see if there are any benefits to choosing B-Splines for the basis functions in KAN layers. Table 14 shows the Top-1 accuracies from different variants of a Deep KAN. In this set of experiments, we closely followed earlier works [5, 6]. Hyperparameters involved in these experiments include the order of the B-spline $k$, grid size $G$, grid range [I,I][-I,I], layers, and width. Although the KANs with B-Spline basis yield high accuracies in the small-scale MNIST and Fashion-MNIST datasets, they fail to generalize over larger datasets (CIFAR-10 and -100).
@@ C.4.1 Experiments with KANs Operated on B-Spline Basis
C.4.2 Ablation with the Hidden Dimension, $r$ in Fourier KArAt
@@
C.4.3 Enforcing an Extreme Low Rank Structure in Fourier KArAt
[…]
C.4.4 Ablation on the Impact of grid size, $G$ on Fourier KArAt
@@
Figure 10: Optimization path for ViT-Tiny and ViT-Base (the smallest and the largest model) along the two largest principal component directions of the successive change of model parameters. We show the loss contours along with the trajectory of the optimizers. Perturbed contours indicate corresponding non-smooth, spiky loss surfaces.
@@
Table 19: Ablation study on the operator variants in Fourier KArAt. Note that $\widehat{\Phi}(W{\cal A})$} and $W\widehat{\Phi}({\cal A})$} refer to order of using linear projector $W$ and operator $\widehat{\Phi}$, and $\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ refers to two consecutive learnable operators $\widehat{\Phi}_{[1]}$ and $\widehat{\Phi}_{[2]}$ applied to $\cal A$}.
@@
C.4.5 Does Fourier KArAt Require $\ell_{1}$-Projection?
@@
C.4.6 More Discussion on Blockwise and Universal Configuration
Although the primary inspiration behind the universal configuration comes from the shared basis functions in KAT [3], it aligns with the fact that softmax is a fixed function conventionally being used across all heads of all encoder layers in the ViT architectures. This raises the question of whether the token-to-token interaction in the self-attention can be modeled using a single universal function, or each instance of attention operations needs dynamic modeling using learnable functions. To investigate further, we compared the learned coefficients $\{a_{pqm}$, $b_{pqm}\}$ from Equation 2 and the weight matrix $W$ of the linear projector following the learnable basis, for each head across all the encoder layers in ViT-Tiny+Fourier KArAt. Thus, we obtain 36 sets (3 heads for each of the 12 layers) of such coefficients and weights from the model with blockwise configuration $G_{3}B$, and 3 sets from the model with universal configuration $G_{3}U$. Although mostly centered at zero, the distribution of each of these sets of parameters varies significantly from the others. The difference is mostly visible in the presence of long tails and the length of the tails. This characteristic is present not only in the blockwise configuration, but the three heads of the universal configuration also show a similar pattern.
C.5 Alternate Variants of Fourier KArAt
C.5.1 Alternative Approaches to the Lower Rank Attention Structures
[…]
C.5.2 Fourier KArAt + softmax Attention —A Hybrid version
C.5.3 More Combinations
C.6 Distribution of the Weights
[74] considered an invariant measure perspective to describe the training loss stabilization of the neural networks. We adopt this idea to study the distribution of the weights of the smallest and largest models of the ViTs, Tiny and Base, and their Fourier KArAt variants. Figure 12 shows the distribution of the weights of these models during different training phases—The evolution of the weights’ distributions for respective models remains invariant. Based on this observation, from [74], we can guarantee the convergence of loss values of all models; also, see Figure 6. However, as mentioned in [74], this perspective cannot comment on the generalization capacity and structural differences between different neural networks.
(a) ViT-Tiny
(b) ViT-Tiny+Fourier KArAt
(c) ViT-Base
(d) ViT-Base+Fourier KArAt
C.7 Spectral Analysis of Attention
We use all 3 heads in the last encoder block of ViT-Tiny on 5 randomly sampled images from the CIFAR-10 validation set. There are a total of 15 singular vectors (each of 197 dimensions) for any attention matrix of shape $197\times 197$, where the singular values are arranged in non-increasing order. Let $\sigma_{i}$ be the $i^{\rm th}$ singular value across all heads and samples (it is permutation invariant). For each $i\in[197],$ we plot $\ln(\sigma_{i}^{\min}),\ln(\bar{\sigma_{i}}),\ln(\sigma_{i}^{\max})]$, where $\bar{\sigma_{i}}$ is the average of $i^{\rm th}$-indexed singular value across all samples and heads.